# Testing Noise Assumptions of Learning Algorithms

### v1

@@ 3 Lower Bounds in the High-Noise Regime
The hard distributions for learning under adversarial label noise proposed by [DKPZ21, Tie23, DKR23] are all indistinguishable from the distribution generated by the oracle $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},1/2}$. Using this fact, we obtain our lower bounds by the following simple observation that any tester-learner that accepts $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},1/2}$ can distinguish between $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},1/2}$ and any distribution where the value of $\mathsf{opt}$ is non-trivial.
@@
  • If $\mathcal{A}$ outputs $\mathsf{Reject}$, then $\mathcal{A}^{\prime}$ outputs 00.
  • If $\mathcal{A}$ outputs $(\mathsf{Accept},h)$, then $\mathcal{A}^{\prime}$ estimates the quantity $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\mathcal{D}_{\mathbf{x},y}}[y\neq h(\mathbf{x})]$ up to tolerance $\tau$ and probability of failure $1/6$ and outputs $1$ if the estimate $q$ is at least $1/2-\tau$ and 00 otherwise.
@@
In the case that $\mathcal{D}_{\mathbf{x},y}\in\mathfrak{D}_{g}$, $\mathcal{A}^{\prime}$ will output 00 unless the guarantee of the soundness does not hold (which happens with probability at most $1/6$) or the error of estimation of $q$ is more than $\tau$ (which happens with probability at most $1/6$. Hence, overall, $\mathcal{A}^{\prime}$ will output $1$ with probability at most $1/3$. ∎
@@ 3.1 Cryptographic Hardness
We provide cryptographic lower bounds based on the widely-believed hardness of the problem of learning with errors (LWE), which was introduced by [Reg09] and is defined as follows.
@@
We receive $m$ i.i.d. examples from some distribution $\mathcal{D}_{\mathbf{x},y}$ over $\mathbb{Z}_{q}^{d}\times\mathbb{R}$ which is either equal to $\mathcal{D}_{\mathsf{null}}$ or $\mathcal{D}_{\mathsf{alt}}$ and we are asked to output $v\in\{\pm 1\}$ such that $|\operatorname*{\mathbb{P}}[v=1|\mathcal{D}_{\mathbf{x},y}=\mathcal{D}_{\mathsf{null}}]-\operatorname*{\mathbb{P}}[v=1|\mathcal{D}_{\mathbf{x},y}=\mathcal{D}_{\mathsf{alt}}]|\geq\alpha$.
There is strong evidence that the LWE problem cannot be solved in subexponential time, since there are quantum reductions from worst-case lattice problems [Reg09, Pei09].
@@
As an immediate corollary of results in [DKR23] (combined with 3.1), we obtain the following lower bound under 3.3.
Theorem 3.4 (Cryptographic Hardness in High-Noise Regime, Theorem 3.1 in [DKR23]).
@@ 3.2 SQ Lower Bounds
We also give lower bounds in the statistical query (SQ) model, which was originally defined by [Kea98]. The SQ framework captures most of the usual algorithmic techniques like moment methods and gradient descent ([FGR+17, FGV17]), and there is a long line of works in computational learning theory giving SQ lower bounds for various learning tasks.
@@
We obtain our lower bound as an immediate corollary of results in [DKR23], combined with 3.1, where note that the reduction of the hard distinguishing problem to testable learning also works in the SQ framework, using one statistical query with tolerance $\tau$.
Theorem 3.6 (SQ Lower Bound in High-Noise Regime, Propositions 2.1, 2.8, Corollary B.1 in [DKPZ21]).
@@ Appendix A Preliminaries
When we say $a=b\pm c$ we mean that $a$ in in the interval $[b-c,b+c]$. When we say a polynomial is “degree-$k$” we mean that the degree polynomial is at most $k$. For a vector $\mathbf{x}$ in $\mathbb{R}^{d}$, let $\mathbf{x}^{\otimes k}$ denote the $\binom{d+1}{k}$-dimensional vector whose elements are of the form $\mathbf{x}^{\alpha}=\prod_{j\in[d]}x_{j}^{\alpha_{j}}$ for $\alpha\in\mathbb{N}^{d}$ with $\|\alpha\|_{1}=\sum_{j}\alpha_{j}\leq k$. In other words, $\mathbf{x}^{\otimes k}$ is the vector one gets by evaluating all multidimensional monomials of degree at most $k$ on input $\mathbf{x}$. We also view degree-$k$ polynomials as corresponding to elements $\mathbb{R}^{\binom{d+1}{k}}$, i.e. their coefficient vectors. Using this notation we have
@@
[…]
To bound the right side of Equation A.3, consider the class $\mathcal{G}$ of $\{0,1\}$-valued functions of the form $\mathbbm{1}_{z\leq p(\mathbf{x})\ \leq R\land\mathbf{x}\in A}$, where $A$ is a set in $\mathcal{H}$ and $p$ is a polynomial in $d$ dimensions of degree at most $d$. Recall that the VC dimension of degree-$k$ polynomial threshold functions is at most $(d+1)^{k}$. From the Sauer-Shelah lemma, it follows that the VC dimension of $\mathcal{G}$ is at most $10((d+1)^{k}+\Delta_{\text{VC}})$. Combining it with the standard VC bound, we see that with probability at least $41-\delta/$ 4 for all monomials $m$ of degree at most $k$ we have
@@
Let $\delta^{\prime}$ be a real number between 00 and $1$, value of which will be chosen later. The Markov’s inequality implies that with probability at least $1-\delta^{\prime}$ we have
@@
Taking $\delta^{\prime}=\frac{\delta}{4(d+1)^{k}}$ and taking a union bound to insure that Equation A.6 holds for all monomials $m$ of degree at most $k$, we see that with probability at least $21-\delta/$ 2 it is the case that all monomials $m$ of degree at most $k$ and all $A$ in $\mathcal{H}$ it is the case that
@@
Recall again that there are at most $\left(d+1\right)^{k}$ degree-$k$ monomials $m$. This allows us to combine Equation A.8 with the triangle inequality to conclude that with probability at least $21-\delta/$ 2 for every polynomial $p$ of degree at most $k$ with coefficients bounded by $B$ in absolute value and for all $A$ in $\mathcal{H}$ we have
@@
The polynomial $p^{2}$ has a degree of at most $2k$ and each coefficient of $p^{2}$ is bounded by $B^{2}\left(d+1\right)^{2k}$. Therefore, with probability at least $21-\delta/$ 2 for every polynomial $p$ of degree at most $k$ with coefficients bounded by $B$ in absolute value and for all $A$ in $\mathcal{H}$ we have
@@ B.2 Soundness
[…]
@@ B.3 Bounding sandwiching degree of the disagreement region
To prove Proposition B.5, we will need the following result by [DGJ+10], [GOWZ10].
@@
Indeed, if $\frac{\log^{3}k}{\sqrt{k}}\leq\frac{1}{t^{2}}$ then the corollary follows from Fact B.7. So all we need to do is to consider the other case. We see that either $t>1$ or $t<-1$ (since $k\geq 2)$. if $t>1$ we take $p_{\text{down}}(\mathbf{x})=0$ and $p_{\text{up}}(\mathbf{x})=\left(\frac{x}{t}\right)^{2}$. If $t<-1$, we take take $p_{\text{up}}(\mathbf{x})=1$ and $p_{\text{down}}(\mathbf{x})=1-\left(\frac{x}{t}\right)^{2}$. In either case, we see that the polynomials $p_{\text{down}}$ and $p_{\text{up}}$ form a pair of $\left(O\left(\min\left(\frac{\log^{3}k}{\sqrt{k}},\frac{1}{t^{2}}\right)\right),1\right)$-sandwiching polynomials of degree $2$. ∎
@@
$\psi_{\text{up}}(\mathbf{x})=\begin{cases}1&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$\geq k_{1}\epsilon$}\text{ or }\mathbf{v}\cdot\mathbf{x}=0\\ \mathbbm{1}_{\mathbf{v}_{\perp}\cdot\mathbf{x}\tan\theta\leq-j\epsilon}&\text{\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$\neq 0$ and } }\mathbf{v}\cdot\mathbf{x}\in\left[j\epsilon,(j+1)\epsilon\right)\text{ for }0\leq j\leq k_{1}-1\\ 0&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$<0$}\end{cases}$
$\psi_{\text{down}}(\mathbf{x})=\begin{cases}0&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$\geq k_{1}\epsilon$}\text{ or }\mathbf{v}\cdot\mathbf{x}=0\\ \mathbbm{1}_{\mathbf{v}_{\perp}\cdot\mathbf{x}\tan\theta<-(j+1)\epsilon}&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$\neq 0$ and }\mathbf{v}\cdot\mathbf{x}\in\left[j\epsilon,(j+1)\epsilon\right)\text{ for }0\leq j\leq k_{1}-1\\ 0&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$<0$}\end{cases}$
@@ B.4 Miscellaneous Claims
Firstly, we compute an orthonormal basis $\left\{\mathbf{w}_{1},\cdots,\mathbf{w}_{d-1}\right\}$ for the $(d-1)$-dimensional subspace of $\mathbb{R}^{d}$ that is orthogonal to $\mathbf{v}$. We express $m(\mathbf{x})=p(\mathbf{w_{1}}\cdot\mathbf{x},\cdots,\mathbf{w_{d-1}}\cdot\mathbf{x},\mathbf{v}\cdot\mathbf{x}$), and note that the polynomial $p$ has all its coefficients between 00 and $\left(d+1\right)^{k_{2}}$, and $p$ is comprised of at most $\left(d+1\right)^{k_{2}}$ monomials. Thus, to have an additive $\beta$-approximation for $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}\left[m(\mathbf{x})\cdot\mathbbm{1}_{a\leq\mathbf{x}\cdot\mathbf{v}<b}\right]$, it sufficies to compute for every monomial $m^{\prime}$ of degree at most $k_{2}$ an additive $\frac{\beta}{d^{2k_{2}}}$-approximation to the quantity
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 3 Lower Bounds in the High-Noise Regime
The hard distributions for learning under adversarial label noise proposed by [diakonikolas2021optimality, tiegel2023hardness, diakonikolas2023near] are all indistinguishable from the distribution generated by the oracle $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},1/2}$. Using this fact, we obtain our lower bounds by the following simple observation that any tester-learner that accepts $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},1/2}$ can distinguish between $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},1/2}$ and any distribution where the value of $\mathsf{opt}$ is non-trivial.
@@
  • If $\mathcal{A}$ outputs $\mathsf{Reject}$, then $\mathcal{A}^{\prime}$ outputs 0.
  • If $\mathcal{A}$ outputs $(\mathsf{Accept},h)$, then $\mathcal{A}^{\prime}$ estimates the quantity $q=\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\mathcal{D}_{\mathbf{x},y}}[y\neq h(\mathbf{x})]$ up to tolerance $\tau$ and probability of failure $1/6$ and outputs $1$ if the estimate $q$ is at least $1/2-\tau$ and 0 otherwise.
@@
In the case that $\mathcal{D}_{\mathbf{x},y}\in\mathfrak{D}_{g}$, $\mathcal{A}^{\prime}$ will output 0 unless the guarantee of the soundness does not hold (which happens with probability at most $1/6$) or the error of estimation of $q$ is more than $\tau$ (which happens with probability at most $1/6$. Hence, overall, $\mathcal{A}^{\prime}$ will output $1$ with probability at most $1/3$. ∎
@@ 3.1 Cryptographic Hardness
We provide cryptographic lower bounds based on the widely-believed hardness of the problem of learning with errors (LWE), which was introduced by [regev2009lattices] and is defined as follows.
@@
We receive $m$ i.i.d. examples from some distribution $\mathcal{D}_{\mathbf{x},y}$ over $\mathbb{Z}_{q}^{d}\times\mathbb{R}$ which is either equal to $\mathcal{D}_{\mathsf{null}}$ or $\mathcal{D}_{\mathsf{alt}}$ and we are asked to output $v\in\{\pm 1\}$ such that $\operatorname*{\mathbb{P}}[v=1|\mathcal{D}_{\mathbf{x},y}=\mathcal{D}_{\mathsf{null}}]-\operatorname*{\mathbb{P}}[v=1|\mathcal{D}_{\mathbf{x},y}=\mathcal{D}_{\mathsf{alt}}]|\geq\alpha$.
There is strong evidence that the LWE problem cannot be solved in subexponential time, since there are quantum reductions from worst-case lattice problems [regev2009lattices, peikert2009public].
@@
As an immediate corollary of results in [diakonikolas2023near] (combined with 3.1), we obtain the following lower bound under 3.3.
Theorem 3.4 (Cryptographic Hardness in High-Noise Regime, Theorem 3.1 in [diakonikolas2023near]).
@@ 3.2 SQ Lower Bounds
We also give lower bounds in the statistical query (SQ) model, which was originally defined by [kearns1998efficient]. The SQ framework captures most of the usual algorithmic techniques like moment methods and gradient descent ([feldman2017statistical, feldman2017statisticalquery]), and there is a long line of works in computational learning theory giving SQ lower bounds for various learning tasks.
@@
We obtain our lower bound as an immediate corollary of results in [diakonikolas2023near], combined with 3.1, where note that the reduction of the hard distinguishing problem to testable learning also works in the SQ framework, using one statistical query with tolerance $\tau$.
Theorem 3.6 (SQ Lower Bound in High-Noise Regime, Propositions 2.1, 2.8, Corollary B.1 in [diakonikolas2021optimality]).
@@ Appendix A Preliminaries
When we say $a=b\pm c$ we mean that $a$ in in the interval [bc,b+c][b-c,b+c]. When we say a polynomial is “degree-kk” we mean that the degree polynomial is at most $k$. For a vector $\mathbf{x}$ in $\mathbb{R}^{d}$, let $\mathbf{x}^{\otimes k}$ denote the $\binom{d+1}{k}-$dimensional vector whose elements are of the form $\mathbf{x}^{\alpha}=\prod_{j\in[d]}x_{j}^{\alpha_{j}}$ for $\alpha\in\mathbb{N}^{d}$ with $\|\alpha\|_{1}=\sum_{j}\alpha_{j}\leq k$. In other words, $\mathbf{x}^{\otimes k}$ is the vector one gets by evaluating all multidimensional monomials of degree at most $k$ on input $\mathbf{x}$. We also view degree-$k$ polynomials as corresponding to elements $\mathbb{R}^{\binom{d+1}{k}}$, i.e. their coefficient vectors. Using this notation we have
@@
$\left|\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}\left[m(\mathbf{x})\mathbbm{1}_{\mathbf{x}\in A\land|m(\mathbf{x})|\leq R\land m(\mathbf{x})>0}\right]-\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}\left[m(\mathbf{x})\mathbbm{1}_{\mathbf{x}\in A\land|m(\mathbf{x})|\leq R\land m(\mathbf{x})>0}\right]\right\rvert=\\ \left|\int_{0}^{R}\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S}\left[m(\mathbf{x})\mathbbm{1}_{\mathbf{x}\in A\land|m(\mathbf{x})|\leq R}\geq z\right]\,dz-\int_{0}^{R}\operatorname*{\mathbb{P}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}\left[m(\mathbf{x})\mathbbm{1}_{\mathbf{x}\in A\land|m(\mathbf{x})|\leq R}\geq z\right]\,dz\right\rvert\leq\\ \int_{0}^{R}\left\lvert\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S}\left[m(\mathbf{x})\mathbbm{1}_{\mathbf{x}\in A\land|m(\mathbf{x})|\leq R}\geq z\right]-\operatorname*{\mathbb{P}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}\left[m(\mathbf{x})\mathbbm{1}_{\mathbf{x}\in A\land|m(\mathbf{x})|\leq R}\geq z\right]\right\rvert\,dz\leq\\ R\max_{z\in[0,R]}\left\lvert\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S}\left[z\leq m(\mathbf{x})\ \leq R\land\mathbf{x}\in A\right]-\operatorname*{\mathbb{P}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}\left[z\leq m(\mathbf{x})\ \leq R\land\mathbf{x}\in A\right]\right\rvert$ (A.3)
To bound the right side of Equation A.3, consider the class $\mathcal{G}$ of $\{0,1\}$-valued functions of the form $\mathbbm{1}_{z\leq p(\mathbf{x})\ \leq R\land\mathbf{x}\in A}$, where $A$ is a set in $\mathcal{H}$ and $p$ is a polynomial in $d$ dimensions of degree at most $d$. Recall that the VC dimension of degree-$k$ polynomial threshold functions is at most $(d+1)^{k}$. From the Sauer-Shelah lemma, it follows that the VC dimension of $\mathcal{G}$ is at most $10((d+1)^{k}+\Delta_{\text{VC}})$. Combining it with the standard VC bound, we see that with probability at least $1-\delta/4$ for all monomials $m$ of degree at most $k$ we have
@@
Let $\delta^{\prime}$ be a real number between 0 and $1$, value of which will be chosen later. The Markov’s inequality implies that with probability at least $1-\delta^{\prime}$ we have
@@
Taking $\delta^{\prime}=\frac{\delta}{4(d+1)^{k}}$ and taking a union bound to insure that Equation A.6 holds for all monomials $m$ of degree at most $k$, we see that with probability at least $1-\delta/2$ it is the case that all monomials $m$ of degree at most $k$ and all $A$ in $\mathcal{H}$ it is the case that
@@
Recall again that there are at most $\left(d+1\right)^{k}$ degree-$k$ monomials $m$. This allows us to combine Equation A.8 with the triangle inequality to conclude that with probability at least $1-\delta/2$ for every polynomial $p$ of degree at most $k$ with coefficients bounded by $B$ in absolute value and for all $A$ in $\mathcal{H}$ we have
@@
The polynomial $p^{2}$ has a degree of at most $2k$ and each coefficient of $p^{2}$ is bounded by $B^{2}\left(d+1\right)^{2k}$. Therefore, with probability at least $1-\delta/2$ for every polynomial $p$ of degree at most $k$ with coefficients bounded by $B$ in absolute value and for all $A$ in $\mathcal{H}$ we have
@@ B.2 Soundness
We say that a function $f:\mathbb{R}^{d}\rightarrow\left\{0,1\right\}$ has $(\epsilon,B)$-sandwiching degree of at most $k$ in $L_{1}$ norm under $\mathcal{N}(0,I_{d})$ with respect to a partition $\mathcal{C}$ of $\mathbb{R}^{d}$ if the function $f$ is $\epsilon$-sandwiched in $L_{1}$ norm under $\mathcal{N}(0,I_{d})$ between $\sum_{A\in\mathcal{C}}\left(p_{\text{down}}^{A}\mathbbm{1}_{A}\right)$ and $\sum_{A\in\mathcal{C}}\left(p_{\text{up}}^{A}\mathbbm{1}_{A}\right)$, where $p_{\text{up}}^{A}$ and $p_{\text{down}}^{A}$ are degreek-k polynomials over $\mathbb{R}^{d}$whose coefficients are bounded by $B$ in absolute value.
@@ B.3 Bounding sandwiching degree of the disagreement region
To prove Proposition B.5, we will need the following result by [diakonikolas2010bounded], [gopalan2010fooling].
@@
Indeed, if $\frac{\log^{3}k}{\sqrt{k}}\leq\frac{1}{t^{2}}$ then the corollary follows from Fact B.7. So all we need to do is to consider the other case. We see that either $t>1$ or t<1t<-1 (since $k\geq 2)$. if $t>1$ we take $p_{\text{down}}(\mathbf{x})=0$ and $p_{\text{up}}(\mathbf{x})=\left(\frac{x}{t}\right)^{2}$. If t<1t<-1, we take take $p_{\text{up}}(\mathbf{x})=1$ and $p_{\text{down}}(\mathbf{x})=1-\left(\frac{x}{t}\right)^{2}$. In either case, we see that the polynomials $p_{\text{down}}$ and $p_{\text{up}}$ form a pair of $\left(O\left(\min\left(\frac{\log^{3}k}{\sqrt{k}},\frac{1}{t^{2}}\right)\right),1\right)$-sandwiching polynomials of degree $2$. ∎
@@
[…]
@@ B.4 Miscellaneous Claims
Firstly, we compute an orthonormal basis $\left\{\mathbf{w}_{1},\cdots,\mathbf{w}_{d-1}\right\}$ for the (d1)(d-1)-dimensional subspace of $\mathbb{R}^{d}$ that is orthogonal to $\mathbf{v}$. We express $m(\mathbf{x})=p(\mathbf{w_{1}}\cdot\mathbf{x},\cdots,\mathbf{w_{d-1}}\cdot\mathbf{x},\mathbf{v}\cdot\mathbf{x}$), and note that the polynomial $p$ has all its coefficients between 0 and $\left(d+1\right)^{k_{2}}$, and $p$ is comprised of at most $\left(d+1\right)^{k_{2}}$ monomials. Thus, to have an additive $\beta$-approximation for $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}\left[m(\mathbf{x})\cdot\mathbbm{1}_{a\leq\mathbf{x}\cdot\mathbf{v}<b}\right]$, it sufficies to compute for every monomial $m^{\prime}$ of degree at most $k_{2}$ an additive $\frac{\beta}{d^{2k_{2}}}$-approximation to the quantity
@@
$\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}\left[m^{\prime}(\mathbf{x})\cdot\mathbbm{1}_{a\leq x_{1}<b}\right]=\underbrace{\left(\prod_{i>1}(\alpha_{i}-1)!!\mathbbm{1}_{\alpha_{i}\text{ is even}}\right)}_{\leq k_{2}^{10k_{2}}}\frac{1}{\sqrt{2\pi}}\int_{a}^{b}e^{-z^{2}/2}z^{\alpha_{1}}\,dz.$
Note that $\alpha_{1}$ is an integer between 0 and $k_{2}$. Since we were seeking to compute a $\frac{\beta}{d^{2k_{2}}}$-approximation to $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}\left[m^{\prime}(\mathbf{x})\cdot\mathbbm{1}_{a\leq x_{1}<b}\right]$, we see that this approximation can be obtained from the equaiton above together with an additive $\frac{\beta}{\left(d+1\right)^{2k_{2}}k^{10k_{2}}}$-approximation to $\frac{1}{\sqrt{2\pi}}\int_{a}^{b}e^{-z^{2}/2}z^{\alpha_{1}}\,dz$. We denote $\rho(z)=e^{-z^{2}/2}z^{\alpha_{1}}$, and let $\beta^{\prime}=\frac{\beta}{\left(d+1\right)^{2k_{2}}k_{2}^{10k_{2}}}$ . We see that the function $\rho$ has the following key properties:
@@
    For all $z_{0}$ in $\mathbb{R}^{d}$ satisfying $z_{0}>4k_{2}+2$ the value $\int_{\left|z\right|>z_{0}}\left|e^{-z^{2}/2}z^{\alpha_{1}}\right|\,dz$ is at most $\int_{\left|z\right|>z_{0}}e^{-z^{2}/4}\,dz$ which in turn is at most $e^{-z_{0}^{2}/4}$.
The three properties above imply that one can approximate the value of $\int_{a}^{b}\rho(z)\,dz$ up to an additive error of $\beta^{\prime}$ via discretization, i.e., by splitting the interval $a,b]\cap[-\sqrt{2\ln(\beta^{\prime})},\sqrt{2\ln\left(\beta^{\prime}\right)}]$ into intervals of size at most $\Delta$ and for each of these intervals $a^{\prime}_{j},b^{\prime}_{j}]$ use the inequality
$\int_{a_{j}^{\prime}}^{b_{j}^{\prime}}\rho(z)\,dz=\rho(a_{j}^{\prime})(a_{j}^{\prime}-b_{j}^{\prime})\pm\left(\sup_{z\in\mathbb{R}}\left|\rho^{\prime}(z)\right|\right)(a_{j}^{\prime}-b_{j}^{\prime})^{2},$
@@
$\int_{z\in[a,b]}\rho(z)\,dz$ $=\overbrace{\int_{z\in[a,b]\cap\left[-\sqrt{2\ln\left(\beta^{\prime}\right)},\sqrt{2\ln\left(\beta^{\prime}\right)}\right]}\rho(z)\,dz\pm\frac{\beta^{\prime}}{2}}^{\text{by property (2) of }\rho}$